The ratio of the number of boys to the number of girls in a room is 5:7. When 24 more boys enter, the ratio is reversed. How many girls are in the room?
a) 29 b) 32 c) 35 d) 36 e) none of these

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

25 boys 35 girls in original room
5*7=35
5/7
10/14
15/21
20/28
25/35
25/35=5/7
(25+24)/35=49/35=7/5
answer 35 girls

Answer by stanbon(75887)

The ratio of the number of boys to the number of girls in a room is 5:7.
5:7 is the same as 5x:7x where 5x is the # of boys
and 7x is the number of girls.
---------------------------------
When 24 more boys enter, the ratio is reversed.
# of boys = 5x+24
# of girls is still 7x
Ratio is 7:5
---
Equation:
(5x+24)/(7x) = 7/5
Cross-multiply:
25x+120 = 49x
24x = 120
x = 5
---
How many girls are in the room?
7x = 35 (# of girls in the room)
Cheers,
Stan H.
a) 29 b) 32 c) 35 d) 36 e) none of these